OpenStudy (anonymous):

what is the slope of a line that is perpendicular to the line represented by 3x+4y=12

OpenStudy (anonymous):

1. 4/3 2. 3/4 3. -3/4 4. -4/3

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (midhun.madhu1987):

3x + 4y = 12 4y = -3x + 12 y = -3/4x + 12/3 y = -3/4x + 4 This is of the form y = mx + c where slope = m = -4/3

OpenStudy (midhun.madhu1987):

where slope = m = -3/4

OpenStudy (anonymous):

no its not that

OpenStudy (midhun.madhu1987):

now product of the slopes of 2 perpendicular line is -1

OpenStudy (midhun.madhu1987):

if one slope is -3/4 then what is the other?

OpenStudy (anonymous):

3/4

OpenStudy (midhun.madhu1987):

nope... try again

OpenStudy (anonymous):

I assume -4/3

OpenStudy (midhun.madhu1987):

its 4/3 because -3/4 * 4/3 = -1

OpenStudy (anonymous):

is it because a negative ?

OpenStudy (anonymous):

ohhhhhhh.

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

think i could have some help on some more?